OpenStudy (jmartinez638):

Calculate the following series:

OpenStudy (jmartinez638):

\[A) \sum_{k=1}^{20}3k^2 + 5k\]

Parth (parthkohli):

\[= 3\sum_{k=1}^{20}k^2 + 5 \sum_{k=1}^{20}k\]Right?

OpenStudy (jmartinez638):

\[B) \lim_{n \rightarrow \infty}\sum_{k=1}^{n}(4 + \frac{ 2k }{ 3 }\times \frac{ 1 }{ n })(\frac{ 1 }{ n })\]

OpenStudy (jmartinez638):

Yes.

Parth (parthkohli):

Now use the results:\[\sum_{k=1}^n k = \frac{n(n+1)}{2}\]\[\sum_{k=1}^n k^2 = \frac{n(n+1)(2n+1)}{6}\]

OpenStudy (jmartinez638):

3(8610)+5(210)

OpenStudy (jmartinez638):

= 26880

Parth (parthkohli):

I guess so.

Parth (parthkohli):

For the second one, just simplify the whole thing and then see what happens when \(n \to \infty\) afterwards

OpenStudy (jmartinez638):

So take the limit after doing the same process to the series?

Parth (parthkohli):

Yeah, don't worry about the limit. We'll handle that later. Just simplify the expression first.

OpenStudy (jmartinez638):

First \[\lim_{n \rightarrow \infty}\sum_{k=1}^{20} (\frac{ 4 }{ n }+\frac{ 2k^3 }{ n }×\frac{ 1 }{ n^2 })\]

Parth (parthkohli):

Where did the 20 come in the picture?

OpenStudy (jmartinez638):

Actually I don't even need to do that. \[4\sum_{k=1}^{n}\frac{ 1 }{ n } + 2\sum_{k=1}^{n}\frac{ k^3 }{ n } + \sum_{k=1}^{n}\frac{ 1 }{ n^2 }\]

OpenStudy (jmartinez638):

Sorry, my mistake

Parth (parthkohli):

Yes, also note that \(n\) is a constant in the summation.

OpenStudy (jmartinez638):

Noted.

Parth (parthkohli):

lol ok, now calculate the rest. don't think about the limit until you're done.

OpenStudy (jmartinez638):

Ok. ((n^2+n)/2) + ((4n^4 + 12n^2)/4) + ((2n^3 + n^2 +3n+n)/6)

OpenStudy (jmartinez638):

Sorry that took so long.

OpenStudy (jmartinez638):

Simplified that would be:

Parth (parthkohli):

oh what

OpenStudy (jmartinez638):

\[\frac{ n^2 + n}{ 2 }+\frac{ 4n^4 +12n^2}{ 4 }+\frac{ 2n^3 +n^2+4n}{ 6 } \rightarrow \frac{ 6n^2 + 6n }{ 12 }+ \frac{ 12n^4 +36n^2}{ 12 } + \frac{ 4n^3 + 2n^2+8n}{ 12 }\]

Parth (parthkohli):

my god no

OpenStudy (jmartinez638):

lol

Parth (parthkohli):

\[\frac{4}{n}+\frac{2k}{3n^2}\]is this your summand?

OpenStudy (jmartinez638):

Sure

OpenStudy (anonymous):

Did anyone calculate the sum to be 9660 ?

OpenStudy (precal):

yes 9660 is correct @robtobey